# PnLCalib: Sports Field Registration via Points and Lines Optimization

### v4

@@ Abstract
Camera calibration in broadcast sports videos presents numerous challenges for accurate sports field registration due to multiple camera angles, varying camera parameters, and frequent occlusions of the field. Traditional search-based methods depend on initial camera pose estimates, which can struggle in non-standard positions and dynamic environments. In response, we propose an optimization-based calibration pipeline that leverages a 3D soccer field model and a predefined set of keypoints to overcome these limitations. Our method also introduces a novel refinement module that improves initial calibration by using detected field lines in a non-linear optimization process. This approach outperforms existing techniques in both multi-view and single-view 3D camera calibration tasks, while maintaining competitive performance in homography estimation. Extensive experimentation on real-world soccer datasets, including SoccerNet-Calibration, WorldCup 2014, and TS-WorldCup, highlights the robustness and accuracy of our method across diverse broadcast scenarios. Our approach offers significant improvements in camera calibration precision and reliability.  [^1] https://github.com/mguti97/PnLCalib
Index Terms:
Camera Calibration, Homography Estimation, Sports Analytics, SoccerNet, World Cup.
@@ II Related work
Sports field registration is a critical component of most sports applications in computer vision, whose common approaches intend to estimate homography matrices in team sports. Traditionally, homography estimation has relied on identifying corresponding features or keypoints between images and the court field model. These features, typically obtained by exploiting geometric primitives such as lines and/or circles, are subsequently used to estimate the mapping between the images. This is often done using the RANSAC algorithm [15] in conjunction with DLT [22] or non-linear optimization techniques [47] that minimize a particular loss function. More recent approaches have diverged from this traditional method. Some directly predict an initial homography matrix, while others seek the optimal matching homography within a reference database containing synthetic images with known homography matrices. Furthermore, the latest approaches have shifted towards directly retrieving camera parameters instead of the homography matrix. This is achieved through various means, including direct prediction, optimization techniques, leveraging databases of image-pose pairs, or decomposing the homography matrix.
@@ II-A Search-based Methods
A prevalent approach in the field has been generating synthetic data to populate databases with homographies or camera poses paired with corresponding image features. These features are often derived from edge maps or semantically segmented images, which represent key elements of the sports field such as lines, circles, and other distinctive markings. Sharma et al. [39] created a synthetic database of edge map-homography pairs. One of the main drawbacks is using normal distribution in the camera pose sampling, which often leads to non-realistic poses. Addressing this issue, Chen and Little [5] created a features-pose database by deducing statistics derived from WorldCup 2014 dataset [23]. Those statistics, related to the camera parameters, were used to sample 90,000 poses which will be encoded through a Siamese Network [20] to distinguish different edge maps. During inference, the network extracts the encoding from edge maps and searches database for nearest neighbour candidate pose. By using area-based semantic segmentation of the soccer field instead of edge images, in contrast with previous approaches, Sha et al. [39] also generated an artificial camera pose database based on the possible ranges for pan and tilt angles and focal length parameters. Aiming to solve image-to-image translation problems, Zhang et al. [52] use improved semantic segmentation using a conditional generative adversarial network [24]. Overall, search-based methods for camera calibration face inherent trade-offs between database size, processing speed, and estimation accuracy. Smaller databases offer faster searches but may compromise initial estimation quality, while larger databases provide better accuracy at the cost of increased computational time. Additionally, these methods often struggle with non-common camera poses frequently encountered in broadcast videos, such as close-ups or oblique angles, which are typically underrepresented in predefined databases.
@@ II-B Optimization-based Methods
Alternatively, common approaches also use edge maps, semantically segmented images or information extracted from the field’s visual landmarks, like intersection points or lines, to obtain homography matrix or camera parameters from optimization methods. Homayounfar et al. [23] classifies field lines using a modified VGG network [41] and extracts their vanishing points reducing the effective number of degrees of freedom (DoF) of the homography from 8 to 4. The field localization problem is formulated as a Markov random field and it requires at least a pair of both vertical and horizontal lines to estimate the vanishing points. Citraro et al. [9] make use of a U-Net architecture [9] to jointly detect semantic keypoints corresponding to field’s line intersections and player positions. In addition to obtaining the homography matrix from the 2D-3D correspondences, intrinsic and extrinsic camera parameters are subsequently obtained through homography decomposition [22]. In order to alleviate the problem of visual landmarks sparsity, Nie et al. [33] propose an encoder-decoder network that jointly outputs a grid of keypoints distributed uniformly on the field template and dense template-features in order to further refine the initial keypoint based homography estimate. Similarly, Maglo et al. [30] use an encoder-decoder network to predict a perspective-aware keypoint grid. This is, the points nearest to the camera are more spread out than the point farthest to the camera in order to compensate for the too large distance variations in the image generated by the perspective effect. More recently, Chu et al. [7] also use a grid of 91 uniformly distributed keypoints and formulate its detection problem as an instance segmentation with dynamic filter learning. This is, the convolution filters are generated dynamically, conditioned on the field image and associated keypoint identity. Oo et al. [34] use a Residual EfficientNet-Attention UNet architecture to estimate the initial homography matrix using pre-defined keypoints to register sports fields. The encoder uses the EfficientNetV2 [44] as the backbone network, and the decoder consists of deconvolution layers with residual blocks, skip connections and attention gates. Theiner et al. [46] introduce a differentiable objective function that is able to learn the camera pose and focal length from segment correspondences. Instance segmentation for each visible line or circle segment is achieved with ResNet [6] backbone and then, the segment reprojection error induced by the estimated camera parameters is iteratively minimized with a gradient-based method [1]. Finally, leveraging the temporal consistency from sports videos, Claasen et al. [10] propose a Bayesian framework. Inspired by recent developments in tracking-by-detection methods, this work proposes a dynamics model that explicitly relates image keypoint positions from one frame to the next through two stages: the first stage consists of a linear Kalman filter, which considers the image keypoints the only part of its state vector, and the second stage incorporates the initial homography estimate to an Extended Kalman Filter [25], with the assumption that the relative homography between frames is small. Falaleev et al. [14] significantly increases the number of usable points for calibration by exploiting line-line and line-conic intersections, points on the conics, and other geometric features, followed by a DLT optimization. Although the computational cost of the enumerated approaches is lower, optimizion-based methods are, overall, not as robust and accurate as search-based methods. Moreover, optimization-based methods are heavily dependent on the landmark detection accuracy.
@@ II-C Prediction-based Methods
Prediction-based approaches use Deep Neural Network-based (DNN) models to calibrate the moving camera in sports. Jiang et al. [26] train a DNN that directly regresses a homography parameterization given an input frame. Subsequently, a sports field template is warped according to the initial estimate. A concatenation of the input frame and the warped template is fed into a second DNN that estimates the error of the current warping and the estimated homography is accordingly optimized until convergence. Tarashima et al. [45] propose SFLNet, a CNN single shot regressor that jointly predicts several outputs: a court metric model defined as an 8-dimensional parameter set, which correspond to the homography’s DoF, a court semantic segmentation which divides the sports field’s spatial layout into divide a frame into court, person, and background regions, and a label adjacency, which comprises adjacencies of label pairs in addition to their presence in an input frame, regularizing the model training via exploiting contextual information. Towards the design of an end-to-end approach, Shi et al. [40] propose a self-supervised learning method for homography estimation. This work employs a self-supervised data mining method to train the registration network with an image and its edge map by using an iterative estimation process controlled by a score regression network to measure the registration error. This method is able to obtain competitive results with previous approaches without the need of any labelled data. Recently, Zhang et al. [53] proposed a four-point calibration method. A cGAN is used to generate semantically segmented frames, eliminating foreground objects. Subsequently, a regression network estimates four points from the frames, which will be used to calculate a homography using DLT algorithm, keeping computational cost low. Lastly, extending to the multiple-view nature of the sports broadcast videos, Mavrogiannis et al. [32] propose a camera calibration method based on synthetic poses. They build a training data generation pipeline with separate flows for the main-camera frames and the rest of camera locations. For the first case, an EfficientNet [6] model is trained to take synthetic edge images as input and regress the camera location. For the second case, a perspective transformation is calculated from four annotated pairs of corresponding points, and the camera location is obtained through homography decomposition. Moreover, a second model is trained on edge maps produced from the poses for each of the possible camera locations to estimate rotation angles and focal length of the camera. Although yielding competitive results on camera calibration and extending it to multiple views, the location-dependent auxiliary model limits the generalization capability and versatility of the method.
@@ II-D Homography and Calibration Refinement
Homography refinement is a crucial step in camera calibration, aiming to achieve an even more accurate homography estimation and camera calibration, when necessary. Previous works use one or a combination of the following methods to refine the initial estimate: Puwein et al. [37] refine the initial homography by bundle adjustment [47]. Chen et al. [5] use the Lucas-Kanade algorithm [3] to refine the initial homography matrix reducing the distance from every pixel in the testing image to its closest edge pixels from a retrieved edge image. Sha et al. [38] introduce the spatial transformer network (STN) to handle large non-affine transformation. The method stacks the input semantic image and the selected template to feed the STN, which outputs a relative homography that maps the semantic image to the template. Citraro et al. [9] use the detected players’ position on the court’s plane to refine the initial homography estimate. After homography decomposition is performed to extract intrinsic and extrinsic camera parameters, these are further refined with Levenberg–Marquardt [22] algorithm. A common approach [52, 53, 32] is to use enhanced correlation coefficient method [13], which performs image alignment between its inputs and returns the refined homography matrix. Alternatively, Oo et al. [34] train an homography refinement network by applying random perturbations to input images. The network outputs the transformation matrix to be applied to the initial estimate. Lastly, other approaches exploit the temporal consistency between subsequent video frames information, such as dense feature maps [33], homography matrices [33, 10] or intrinsic and/or extrinsic camera parameters [9].
@@ III Methodology
Sports TV broadcasts consist of video sequences featuring a fraction of the sports field from different uncalibrated moving camera perspectives, defining a multiple-view setting. Our approach focuses on retrieving both extrinsic and intrinsic camera parameters from each individual frame, without any prior information about the camera’s position or orientation, except for a partial view of the soccer field. The proposed method comprises five processing components: soccer field modelling and keypoints generation, keypoints and line extremities detection, DLT algorithm and camera parameters retrieval, as well as calibration estimate refinement. Next, these components are introduced.
III-A Modelling the Soccer Field
A soccer field is composed of lines, circles and semi-circles, representing all field markings, goal posts, and crossbars. Our approach, following keypoint-based methods [45, 9, 33, 7, 30, 10, 34, 14], relies on the lines painted on the ground, their intersections, the corners they define and some extra geometric properties, due to its known position on the world coordinate system. We follow the segment definitions of Cioppa et al. [8] and set them as starting points to hierarchically compute our pre-defined keypoint grid using court geometric properties.
@@ III-A1 Keypoint Generation
  • Extended Line-Line intersections. Following [11], this set ($\mathcal{K}p_{e}$) addresses the intersections of extended lines that represent non-adjacent segments of the soccer field. To obtain this set, the field lines are extended by exploiting their line equations beyond their original boundaries. It is worth noting that not all non-adjacent line intersections are added; if the lines have to be largely extended, small errors in the line equation would lead to huge deviations in the intersection position.
@@ III-B Keypoints and Lines Detection
Our approach is built upon Falaleev et al. [14] solution, which makes use of two encoder-decoder convolutional neural networks to estimate the position of the pre-defined keypoints and the soccer field lines depicted in [8] excluding conics. While in [14] the line model is given an auxiliary role to enhance keypoint completeness, in our approach is used as the key component of our refinement module in the last calibration stage. During inference, the former produces heatmaps for each pre-defined keypoint with a single $2$-pixel Gaussian peak sigma positioned in the keypoint location, accompanied by an additional background channel. This additional channel reflects the inverse of the maximum value among the other target feature maps, ensuring that the resultant target tensor behaves as a probability distribution function at every spatial point. Meanwhile, the latter network produces heatmaps for each visible soccer field line within the frame, assigning two Gaussian peaks at the line extremities’ locations. Additionally, we introduce an extra channel, known as the boundary channel, to our heatmap following the approach outlined in [50]. This augmentation aims to enhance the efficient capture of global information regarding the soccer field and improve extremities detection, particularly near image borders. We effectively extract the position of keypoints and line extremities from the generated heatmaps by employing a max pooling operation, which calculates the maximum value for patches of a feature map, drawing inspiration from the methodology proposed in [55]. This process is summarized on Fig. 1-Bottom.

### v5

@@ Abstract
Camera calibration in broadcast sports videos presents numerous challenges for accurate sports field registration due to multiple camera angles, varying camera parameters, and frequent occlusions of the field. Traditional search-based methods depend on initial camera pose estimates, which can struggle in non-standard positions and dynamic environments. In response, we propose an optimization-based calibration pipeline that leverages a 3D soccer field model and a predefined set of keypoints to overcome these limitations. Our method also introduces a novel refinement module that improves initial calibration by using detected field lines in a non-linear optimization process. This approach outperforms existing techniques in both multi-view and single-view 3D camera calibration tasks, while maintaining competitive performance in homography estimation. Extensive experimentation on real-world soccer datasets, including SoccerNet-Calibration, WorldCup 2014, and TS-WorldCup, highlights the robustness and accuracy of our method across diverse broadcast scenarios. Our approach offers significant improvements in camera calibration precision and reliability. Our project is available at https://github.com/mguti97/PnLCalib.
@@ I Introduction
Recent developments in camera calibration have moved beyond traditional handcrafted approaches and now encompass a wide spectrum of learning-based techniques tailored to various camera models and deployment scenarios. For standard pinhole cameras, deep learning methods have demonstrated the ability to estimate both intrinsic and extrinsic parameters from image content alone. Early work like DeepFocal [66] focused on focal length regression, while PoseNet [32] introduced the use of CNNs for full 6-DoF pose estimation. Subsequent methods have incorporated geometric reasoning through intermediate representations such as vanishing points [68] or horizon lines [67], enabling more robust and interpretable estimation pipelines. DeepFEPE [29] and NeurVPS [73], for instance, improve pose estimation by integrating differentiable modules that mimic traditional keypoint detection and geometric constraint solving. These methods benefit from their ability to generalize across scenes without the need for known targets or structured environments. In parallel, the calibration of distortion models has also advanced significantly. For radial and rolling shutter distortion, deep networks have been used to learn corrective transformations directly from image data, using either parameter regression or dense distortion maps. Techniques like URS-CNN [49] and DeepUnrollNet [34] have proven effective for modeling image formation under non-ideal optical conditions. Beyond single-camera settings, cross-view calibration addresses relative pose or homography estimation between images captured from different viewpoints. Approaches such as UDHN [43], BasesHomo [25], and HomoGAN [27] use supervised or unsupervised objectives to learn geometric alignment across diverse perspectives, often without access to ground truth calibration. These are especially useful when image pairs exhibit wide baselines or temporal variation. Cross-sensor calibration further extends the calibration problem to systems involving different modalities—such as RGB cameras and LiDAR. Networks like RegNet [50], CFNet [37], and SemAlign [35] estimate spatial correspondences by aligning semantic, geometric, or point cloud features across sensors. Moreover, calibration setups that leverage external devices—such as collimator systems [33]—offer a practical alternative in environments where target-based calibration is infeasible due to long working distances or spatial constraints. Complementing these advances, auto-calibration methods have continued to evolve, offering solutions that do not rely on dedicated calibration targets or full projective reconstructions. Classical approaches  [61, 5] depend on strong assumptions, such as knowledge of multiple fundamental matrices or fixed internal parameters across views. In contrast, newer methods explore more flexible formulations that make use of partially known intrinsics or multi-view image correspondences. Recent work [9] has further shown that by leveraging minimal geometric constraints.
@@ II Related Work
Sports field registration is a critical component of most sports applications in computer vision, whose common approaches intend to estimate homography matrices in team sports. Traditionally, homography estimation has relied on identifying corresponding features or keypoints between images and the court field model. These features, typically obtained by exploiting geometric primitives such as lines and/or circles, are subsequently used to estimate the mapping between the images. This is often done using the RANSAC algorithm [17] in conjunction with DLT [24] or non-linear optimization techniques [60] that minimize a particular loss function. More recent approaches have diverged from this traditional manner. Some directly predict an initial homography matrix, while others seek the optimal matching homography within a reference database containing synthetic images with known homography matrices. Furthermore, the latest approaches have shifted towards directly retrieving camera parameters instead of the homography matrix. This is achieved through various means, including direct prediction, optimization techniques, leveraging databases of image-pose pairs, or decomposing the homography matrix.
@@ II-A Search-based Methods
A prevalent approach in the field has been generating synthetic data to populate databases with homographies or camera poses paired with corresponding image features. These features are often derived from edge maps or semantically segmented images, which represent key elements of the sports field such as lines, circles, and other distinctive markings. [52] created a synthetic database of edge map-homography pairs. One of the main drawbacks is using normal distribution in the camera pose sampling, which often leads to non-realistic poses. Addressing this issue, [6] created a features-pose database by deducing statistics derived from WorldCup 2014 dataset [26]. Those statistics, related to the camera parameters, were used to sample 90,000 poses which will be encoded through a Siamese Network [22] to distinguish different edge maps. During inference, the network extracts the encoding from edge maps and searches database for nearest neighbor candidate pose. By using area-based semantic segmentation of the soccer field instead of edge images, in contrast with previous approaches, [52] also generated an artificial camera pose database based on the possible ranges for pan and tilt angles and focal length parameters. Aiming to solve image-to-image translation problems, [69] use improved semantic segmentation by exploiting a conditional generative adversarial network [28]. Overall, search-based methods for camera calibration face inherent trade-offs between database size, processing speed, and estimation accuracy. Smaller databases offer faster searches but may compromise initial estimation quality, while larger databases provide better accuracy at the cost of increased computational time. Additionally, these methods often struggle with non-common camera poses frequently encountered in broadcast videos, such as close-ups or oblique angles, which are typically underrepresented in predefined databases.
@@ II-B Optimization-based Methods
Alternatively, common approaches also use edge maps, semantically segmented images or information extracted from the field’s visual landmarks, like intersection points or lines, to obtain homography matrix or camera parameters from optimization methods. [26] classify field lines using a modified VGG network [54] and extracts their vanishing points reducing the effective number of degrees of freedom (DoF) of the homography from 8 to 4. The field localization problem is formulated as a Markov random field and it requires at least a pair of both vertical and horizontal lines to estimate the vanishing points. [11] make use of a U-Net architecture [11] to jointly detect semantic keypoints corresponding to field’s line intersections and player positions. In addition to obtaining the homography matrix from the 2D-3D correspondences, intrinsic and extrinsic camera parameters are subsequently obtained through homography decomposition [24]. In order to alleviate the problem of visual landmarks sparsity, [44] propose an encoder-decoder network that jointly outputs a grid of keypoints distributed uniformly on the field template and dense template-features in order to further refine the initial keypoint based homography estimate. Similarly, [40] use an encoder-decoder network to predict a perspective-aware keypoint grid. This is, the points nearest to the camera are more spread out than the point farthest to the camera in order to compensate for the too large distance variations in the image generated by the perspective effect. More recently, [8] also use a grid of 91 uniformly distributed keypoints and formulate its detection problem as an instance segmentation with dynamic filter learning. This is, the convolution filters are generated dynamically, conditioned on the field image and associated keypoint identity. [45] use a Residual EfficientNet-Attention UNet architecture to estimate the initial homography matrix using pre-defined keypoints to register sports fields. The encoder uses the EfficientNetV2 [57] as the backbone network, and the decoder consists of deconvolution layers with residual blocks, skip connections and attention gates. [59] introduce a differentiable objective function that is able to learn the camera pose and focal length from segment correspondences. Instance segmentation for each visible line or circle segment is achieved with ResNet [7] backbone and then, the segment reprojection error induced by the estimated camera parameters is iteratively minimized with a gradient-based method [1]. Finally, leveraging the temporal consistency from sports videos, [12] propose a Bayesian framework. Inspired by recent developments in tracking-by-detection methods, this work proposes a dynamics model that explicitly relates image keypoint positions from one frame to the next through two stages: the first stage consists of a linear Kalman filter, which considers the image keypoints the only part of its state vector, and the second stage incorporates the initial homography estimate to an Extended Kalman Filter [30], with the assumption that the relative homography between frames is small. [16] significantly increase the number of usable points for calibration by exploiting line-line and line-conic intersections, points on the conics, and other geometric features, followed by a DLT optimization. Although the computational cost of the enumerated approaches is lower, optimization-based methods are, overall, not as robust and accurate as search-based methods. Moreover, optimization-based methods are heavily dependent on the landmark detection accuracy.
@@ II-C Prediction-based Methods
Prediction-based approaches use Deep Neural Network-based (DNN) models to calibrate the moving camera in sports. [31] train a DNN that directly regresses a homography parameterization given an input frame. Subsequently, a sports field template is warped according to the initial estimate. A concatenation of the input frame and the warped template is fed into a second DNN that estimates the error of the current warping and the estimated homography is accordingly optimized until convergence. [58] propose SFLNet, a CNN single shot regressor that jointly predicts several outputs: a court metric model defined as an 8-dimensional parameter set, which correspond to the homography’s DoF, a court semantic segmentation which divides the sports field’s spatial layout into divide a frame into court, person, and background regions, and a label adjacency, which comprises adjacency of label pairs in addition to their presence in an input frame, regularizing the model training via exploiting contextual information. Towards the design of an end-to-end approach, [53] propose a self-supervised learning method for homography estimation. This work employs a self-supervised data mining method to train the registration network with an image and its edge map by using an iterative estimation process controlled by a score regression network to measure the registration error. This method is able to obtain competitive results with previous approaches without the need of any labeled data. Recently, [70] propose a four-point calibration method. A cGAN is used to generate semantically segmented frames, eliminating foreground objects. Subsequently, a regression network estimates four points from the frames, which will be used to calculate a homography using DLT algorithm, keeping computational cost low. Lastly, extending to the multiple-view nature of the sports broadcast videos, [42] propose a camera calibration method based on synthetic poses. They build a training data generation pipeline with separate flows for the main-camera frames and the rest of camera locations. For the first case, an EfficientNet [7] model is trained to take synthetic edge images as input and regress the camera location. For the second case, a perspective transformation is calculated from four annotated pairs of corresponding points, and the camera location is obtained through homography decomposition. Moreover, a second model is trained on edge maps produced from the poses for each of the possible camera locations to estimate rotation angles and focal length of the camera. Although yielding competitive results on camera calibration and extending it to multiple views, the location-dependent auxiliary model limits the generalization capability and versatility of the method.
@@ II-D Homography and Calibration Refinement
Homography refinement is a crucial step in camera calibration, aiming to achieve an even more accurate homography estimation and camera calibration, when necessary. Previous works use one or a combination of the following methods to refine the initial estimate: [48] refine the initial homography by bundle adjustment [60]. [6] use the Lucas-Kanade algorithm [3] to refine the initial homography matrix reducing the distance from every pixel in the testing image to its closest edge pixels from a retrieved edge image. [51] introduce the spatial transformer network (STN) to handle large non-affine transformation. The method stacks the input semantic image and the selected template to feed the STN, which outputs a relative homography that maps the semantic image to the template. [11] use the detected players’ position on the court’s plane to refine the initial homography estimate. After homography decomposition is performed to extract intrinsic and extrinsic camera parameters, these are further refined with Levenberg–Marquardt [24] algorithm. A common approach [69, 70, 42] is to use enhanced correlation coefficient method [15], which performs image alignment between its inputs and returns the refined homography matrix. Alternatively, [45] train an homography refinement network by applying random perturbations to input images. The network outputs the transformation matrix to be applied to the initial estimate. Lastly, other approaches exploit the temporal consistency between subsequent video frames information, such as dense feature maps [44], homography matrices [44, 12] or intrinsic and/or extrinsic camera parameters [11].
@@ III Methodology
Sports TV broadcasts consist of video sequences featuring a fraction of the sports field from different uncalibrated moving camera perspectives, defining a multiple-view setting. Our approach focuses on retrieving both extrinsic and intrinsic camera parameters from each individual frame, without any prior information about the camera’s position or orientation, except for a partial view of the soccer field. The proposed open-source method comprises five processing components: soccer field modeling and keypoints generation, keypoints and line extremities detection, DLT algorithm and camera parameters retrieval, as well as calibration estimate refinement. Next, these components are introduced.
III-A Modeling the Soccer Field
A soccer field is composed of lines, circles and semi-circles, representing all field markings, goal posts, and crossbars. Our approach, following keypoint-based methods [58, 11, 44, 8, 40, 12, 45, 16], relies on the lines painted on the ground, their intersections, the corners they define and some extra geometric properties, due to its known position on the world coordinate system. We follow the segment definitions of [10] and set them as starting points to hierarchically compute our pre-defined keypoint grid using court geometric properties.
@@ III-A1 Keypoint Generation
  • Extended Line-Line intersections. Following [13], this set ($\mathcal{K}p_{e}$) addresses the intersections of extended lines that represent non-adjacent segments of the soccer field. To obtain this set, the field lines are extended by exploiting their line equations beyond their original boundaries. It is worth noting that not all non-adjacent line intersections are added; if the lines have to be largely extended, small errors in the line equation would lead to huge deviations in the intersection position. To address this, only intersections located within the image boundaries—or at most up to 50% of the image size beyond them—are added to the set.
@@ III-B Keypoints and Lines Detection
Our approach is built upon [16] solution, which makes use of two encoder-decoder convolutional neural networks to estimate the position of the pre-defined keypoints and the soccer field lines depicted in [10] excluding conics. While in [16] the line model is given an auxiliary role to enhance keypoint completeness, in our approach is used as the key component of our refinement module in the last calibration stage. During inference, the former produces heatmaps for each pre-defined keypoint with a single $2$-pixel Gaussian peak sigma positioned in the keypoint location, accompanied by an additional background channel. This additional channel reflects the inverse of the maximum value among the other target feature maps, ensuring that the resultant target tensor behaves as a probability distribution function at every spatial point. Meanwhile, the latter network produces heatmaps for each visible soccer field line within the frame, assigning two Gaussian peaks at the line extremities’ locations. Additionally, we introduce an extra channel, known as the boundary channel, to our heatmap following the approach outlined in [64]. This augmentation aims to enhance the efficient capture of global information regarding the soccer field and improve extremities detection, particularly near image borders. We effectively extract the position of keypoints and line extremities from the generated heatmaps by employing a max pooling operation, which calculates the maximum value for patches of a feature map, drawing inspiration from the methodology proposed in [72]. This process is summarized on Fig. 1-Bottom.